Tom’s uncle owns a triangular piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one-half of the longest side. The second longest side is 2 yards shorter than the longest side. What is the length of each side?
What does this problem mean? I tried and I don't understand it. Please help me

Answer by checkley71(8403)

Longest side=L
Shortest side=(l/2*L+30)
Second longest side=(L-2)
L+(L/2+30)+(L-2)=378
L+L/2+30+L-2=378
2.5L=378-30+2
2.5L=350
L=350/2.5
L=140 THE LENGTH OF THE LONGEST SIDE.
140/2+30=70+30=100 LENGTH OF THE SHORTEST SIDE.
140-2=138 THE SECOND LONGEST SIDE.
PROOF:
140+100+138=378
378=378